OpenStudy (anonymous):

How many solutions exist for the system y = 2x + 5 and y = -0.2x - 4? A. no solution B. one solution C. infinitely many solutions

OpenStudy (anonymous):

B

OpenStudy (anonymous):

y = 2x + 5 and y = -0.2x - 4 2x + 5 = -0.2x - 4 2.2x=-9 x=9/2.2

OpenStudy (anonymous):

one solution.